Flex Co. just paid total dividends of $1,125,000 and reported additions to retained earnings of $3,375,000. The company has 735,000 shares of stock outstanding and a benchmark PE of 17.50 times. What stock price would you consider appropriate?

$96.43

$26.79

$80.36

$107.14

$101.79
Answer #1

Toral dividend = 1125000 Retained eaming - 3375000 Mo. of shares = 735000 P/E = 17-50 times. Coming /share= (Dividend & addit